At the beginning of the year, Gonzales Company SLU had total assets of Є844,740 and total liabilities of Є407,570. Answer the following independent questions.

(a) If total assets increased Є165,880 during the year and total liabilities decreased Є61,120, what is the amount of equity at the end of the year?

(b) During the year, total liabilities increased Є103,920 and equity decreased Є62,760. What is the amount of total assets at the end of the year?

(c) If total assets decreased Є66,370 and equity increased Є88,490 during the year, what is the amount of total liabilities at the end of the year?
Answer #1

For Gonzalez Company SLU

At the beginning of the year:

Total Assets = 844,740

Total Liabilities = 407,570

Total Assets = Total Liabilities + Equity

Thus, Equity at the beginning of the year = Total Assets – Total Liabilities = 844,740 – 407,570 = 437,170

a. Increase in Total Assets = 165,880

Total Assets at the end of the year = Total Assets at the beginning of the year + Increase in Total Assets = 844,740 + 165,880 = 1,010,620

Decrease in Total Liabilities = 61,120

Total Liabilities at the end of the year = 407,570 – 61,120 = 346,450

Equity at the end of the year = Total Assets at the end of the year – Total Liabilities at the end of the year = 1,010,620 – 346,450 = 664,170

b. Increase in Total Liabilities = 103,920

Total Liabilities at the end of the year = Total Liabilities at the beginning of the year + Increase in Total Liabilities = 407,570 + 103,920 = 511,490

Decrease in Equity = 62,760

Equity at the end of the year = Equity at the beginning of the year – Decrease in Equity = 437,170 – 62,760 = 374,410

Total Assets at the end of the year = Total Liabilities at the end of the year + Equity at the ed of the year = 511,490 + 374,410 = 885,900

c. Decrease in Total Assets = 66,370

Total Assets at the end of the year = Total Assets at the beginning of the year – Decrease in Total Assets = 844,740 – 66,370 = 778,370

Increase in Equity = 88,490

Equity at the end of the year = Equity at the beginning of the year + Increase in Equity = 437,170 + 88,490 = 525,660

Total Liabilities at the end of the year = Total Assets at the end of the year – Equity at the end of the year = 778,370 – 525,660 = 252,710